# Learned Display Radiance Fields with Lensless Cameras

### v1

@@ Abstract.
Calibrating displays is a basic and regular task that content creators must perform to maintain optimal visual experience, yet it remains a troublesome issue. Measuring display characteristics from different viewpoints often requires specialized equipment and a dark room, making it inaccessible to most users. To avoid specialized hardware requirements in display calibrations, our work co-designs a lensless camera and an Implicit Neural Representation based algorithm for capturing display characteristics from various viewpoints. More specifically, our pipeline enables efficient reconstruction of light fields emitted from a display from a viewing cone of $46.6^{\circ}\times 37.6^{\circ}$. Our emerging pipeline paves the initial steps towards effortless display calibration and characterization.
Lensless camera, Implicit Representation, Display calibration
@@ 1. Introduction
Display chromaticity and luminance vary with viewing angle, affecting how users perceive content from different positions in front of a screen. But most calibration tools assume a fixed viewing position leading to invalid assessments for other viewing positions. This is not only problematic for display manufacturers but also for professionals that need consistent color edits for various geometric configurations. The ISO standards (International Organization for Standardization, 2008) define procedures for assessing the uniformity distribution of Flat Panel Displays (FPD) at various viewing directions by rotating either the display or measurement instrument in a dark room. However, these requirements with numerous captures hinder the end users from performing regular calibration, which is critical for applications such as virtual production display walls  [^4] https://partnerhelp.netflixstudios.com/hc/en-us/articles/1500002086641-Common-Virtual-Production-Challenges-Potential-Solutions, professional workstations in graphic design, video editing  [^5] https://displaycal.net/, and automotive dashboard systems (Kim et al., 2022). An accessible method with a reasonable amount of captures is needed to measure the image quality of the displays from arbitrary viewpoints.
@@
We conduct experiments on a Liquid Crystal display (LCD) display and demonstrate that our method can reconstruct the display with different viewpoints. Our codebase and hardware design will be available upon acceptance.
@@ 2. Methods
Our method reconstructs the light field representation of a display from lensless captures of individual pixels. We detail our pipeline consisting hardware and software components in the subsequent sections. The notations used throughout this section are summarized in Tbl. 1. Upper-case bold letters denote matrices, while lower-case bold letters represent vectors. Scalars and spatial indices are denoted by non-boldface letters. The calligraphic letters indicate the functions. These notations will be consistently used to formalize the forward model and reconstruction process in the following subsections.
@@ 2.1. Lensless 2D Image Formation
where $\mathbf{y}_{I}\in\mathbb{R}^{(h_{I}\cdot w_{I})\times 1}$ denotes the flattened measurement vector from the sensor, $\mathbf{H}\in\mathbb{R}^{(h_{I}\cdot w_{I})\times(h_{S}\cdot w_{S})}$ is a coefficient matrix modeling the input light’s interaction with the diffuser, and $\mathbf{x}_{S}\in\mathbb{R}^{(h_{S}\cdot w_{S})\times 1}$ represents the flattened intensity vector of the unknown scene. However, reconstructing the full $\mathbf{H}$ matrix is computationally intractable. To address this, we position the diffuser close to the sensor to ensure spatial invariance in the forward model, allowing it to be approximated as a linear convolution:
@@
$=\mathcal{F}^{-1}\left(\mathcal{F}\big(\mathcal{C}(\mathcal{Z}(\mathbf{X}))\big)\cdot\mathcal{F}\big(\mathcal{Z}(\mathbf{h})\big),\right)$
@@ 2.2. Lensless Light Field Formation
(5) $\mathbf{I}_{\text{\text{gt}}}=\mathcal{F}^{-1}\Bigg(\mathcal{F}\Big(\mathcal{C}\big(\mathcal{Z}(\mathbf{L})\big)\Big)\cdot\mathcal{F}\Big(\mathcal{Z}(\mathfrak{h})\Big)\Bigg).$
@@
(6) $\hat{\mathbf{L}}\leftarrow\underset{\mathbf{L}}{\mathrm{argmin}}\;\mathcal{L}\Big(\mathbf{I}_{\text{\text{gt}}},\mathcal{C}\big(\mathcal{Z}(L)*\mathcal{Z}(\mathfrak{h})\big)\Big),$
@@ 2.3. Implicit Neural Representation
where $\mathbf{F}_{\theta}$ is an MLP with parameters $\theta$. This network takes the spatial and angular coordinates as input and outputs the corresponding pixel color value. The variables $\mathbf{x}$ and $\mathbf{y}$ denote the illuminated pixel coordinates on the display. Before the input coordinates are fed into the MLP, we apply positional encoding (Mildenhall et al., 2021) to the coordinates to increase the model’s capacity to learn high-frequency details:
@@ 3. Evaluation and Discussion
Our hardware setup consists of a $0.5^{\circ}$ engineered diffuser (Edmund 35-860) with an array of five $2.5\,mm\times 3\,mm$ apertures serving as a mask positioned 10mm10\,$m$ from the imaging sensor as illustrated in Fig. 4. This distance is experimentally determined to ensure spatial invariance of the diffuser and the sharpest caustic pattern, satisfying condition in Eq. (2). We 3D print a housing that maintains a fixed 30mm30\,$m$ distance between the diffuser and screen pixel, as the left image in Fig. 5 demonstrates. We capture a stack of PSF for each aperture by positioning the lensless camera in front of a white LED coupled to a $100\,\mu m$ pinhole (Thorlabs P100HKb). We mount the prototype in front of the display panel and turn on each pixel one by one to capture the light fields. We choose the position of activated pixel relative to the corresponding aperture to ensure the light rays pass through the aperture and reach the imaging sensor. We detail this process in Supplementary Sec. 1.
@@
We divide the lensless captures into $9\times 9$ sub-aperture images using a sliding window of size $54\times 70$ pixels. Later we feed them into our MLP that consists of 3 fully connected layers with sinusoidal activation functions, each containing 32 channels. Before inputting the coordinates described in Sec. 2.3, we inject random noise into each coordinate group, which helps improve generalization: display coordinates (std: $5\times 10^{-3}$), angular coordinates (std: $1\times 10^{-2}$), and subview coordinates (std: $1\times 10^{-3}$). Subsequently, we apply the positional encoding function Eq. (8) with one frequency level for display coordinates, five for angular coordinates, and ten for spatial coordinates, as illustrated in Fig. 3. We analyze the choices of frequency levels in Supplementary Sec. 2 to ensure the model is generalized to unseen pixel position and viewpoints. We optimize our model using the Adam optimizer with an initial learning rate of $0.001$ that decays linearly over $800$ epochs. To prevent gradient explosion during early training, gradient norms are clipped to $1.0$. We train the model on independent color channels to mitigate crosstalk, which takes about one hour on a single NVIDIA RTX 2070 GPU, while inference requires roughly $0.01$ seconds per 1080p frame.
@@
We assess pixel consistency by measuring how well the reconstructed light fields preserve intensity variations across continuous horizontal and vertical incident angles. Fig. 5 shows that the average pixel intensity changes smoothly, indicating that our model can predict novel, angle-dependent pixel values. We compare our architecture against a vanilla MLP baseline consisting of fully connected layers with ReLU activations in Tbl. 2. Our method achieves higher scores in all metrics, demonstrating superior performance.
@@
To benchmark our method, we follow the ISO (International Organization for Standardization, 2008) standard: a conventional camera is aligned with the display center in a dark room, full-screen white stimuli are displayed, and images are captured across vertical incident angles from $10^{\circ}$ to $16^{\circ}$. The average intensity curves in Fig. 6 show that our method reproduces the similar trend in the ISO setup, validating its physical plausibility. Unlike the ISO standard, our lensless camera setup captures light field data from nine positions on the display without rotating the camera or using controlled lighting. This reduces measurement time and simplifies view-dependent calibration, while maintaining accuracy comparable to the ISO method.
@@
Methods PSNR (dB) $\uparrow$ SSIM $\uparrow$ MSSIM $\uparrow$
Ours 19.54 0.9165 0.9549
Vanilla 9.93 0.5327 0.8049
@@
Our $46.6^{\circ}$ angular coverage is below the $240^{\circ}$ specified in professional display calibration. Future improvements with larger aperture arrays and optimized diffusers could extend this range. Additionally, pixel-wise training limits scalability for full-panel characterization. Several research directions offer potential solutions to these limitations. One is an end-to-end pipeline that removes PSF convolutions, cropping, and padding. Another is the use of emerging Gaussian splatting representations, which model light fields with fewer parameters than INR.

### v2

@@ Abstract.
Calibrating displays is a basic and regular task that content creators must perform to maintain optimal visual experience, yet it remains a troublesome issue. Measuring display characteristics from different viewpoints often requires bulky equipment and a dark room, making it inaccessible to most users. To avoid such hardware requirements in display calibrations, our work co-designs a lensless camera and an Implicit Neural Representation based algorithm for capturing display characteristics from various viewpoints. More specifically, our pipeline enables efficient reconstruction of light fields emitted from a display from a viewing cone of $46.6^{\circ}\times 37.6^{\circ}$. Our emerging pipeline paves the initial steps towards effortless display calibration and characterization.
@@ 1. Introduction
Display chromaticity and luminance vary with viewing angle, affecting how users perceive content from different positions in front of a screen. But most calibration tools assume a fixed viewing position leading to invalid assessments for other viewing positions. This is not only problematic for display manufacturers but also for professionals that need consistent color edits for various geometric configurations. The ISO standards (ISO, 2008) define procedures for assessing the uniformity distribution of Flat Panel Displays (FPD) at various viewing directions by rotating either the display or measurement instrument in a dark room. However, these requirements with numerous captures hinder the end users from performing regular calibration, which is critical for virtual production display walls  [^4] https://partnerhelp.netflixstudios.com/hc/en-us/articles/1500002086641-Common-Virtual-Production-Challenges-Potential-Solutions, professional graphic design, and video editing  [^5] https://displaycal.net/. An accessible method with a reasonable amount of captures is needed to measure the image quality of the displays from arbitrary viewpoints.
@@
We conduct experiments on a Liquid Crystal Display (LCD) and demonstrate that our method can reconstruct the display with different viewpoints. Our codebase is available at GitHub:complight.
@@ 2. Methods
Our method reconstructs the light field representation of a display from lensless captures of individual pixels. We detail our pipeline consisting of hardware and software components in the subsequent sections. Our notations are summarized in Tbl. 1. Upper-case bold letters denote matrices, while lower-case bold letters represent vectors. Scalars and spatial indices are denoted by non-boldface letters. The calligraphic letters indicate the functions.
@@ 2.1. Lensless 2D Image Formation
where $\mathbf{y}_{I}\in\mathbb{R}^{(h_{I}\cdot w_{I})}$ denotes the flattened measurement vector from the sensor, $\mathbf{H}\in\mathbb{R}^{(h_{I}\cdot w_{I})\times(h_{S}\cdot w_{S})}$ is a coefficient matrix modeling the input light’s interaction with the diffuser, and $\mathbf{x}_{S}\in\mathbb{R}^{(h_{S}\cdot w_{S})}$ represents the flattened intensity vector of the unknown scene. However, reconstructing the full $\mathbf{H}$ matrix is computationally intractable. To address this, we position the diffuser close to the sensor to ensure spatial invariance in the forward model, allowing it to be approximated as a linear convolution:
@@
$=\mathcal{C}\left(\mathcal{F}^{-1}\left(\mathcal{F}\big(\mathcal{Z}(\mathbf{X})\big)\cdot\mathcal{F}\big(\mathcal{Z}(\mathbf{h})\big)\right)\right),$
@@ 2.2. Lensless Light Field Formation
(5) $\mathbf{I}_{\text{\text{gt}}}=\mathcal{C}\left(\mathcal{F}^{-1}\left(\mathcal{F}\big(\mathcal{Z}(\mathbf{L})\big)\cdot\mathcal{F}\big(\mathcal{Z}(\mathfrak{h})\big)\right)\right).$
@@
(6) $\hat{\mathbf{L}}\leftarrow\underset{\mathbf{L}}{\mathrm{argmin}}\;\mathcal{L}\Big(\mathbf{I}_{\text{\text{gt}}},\mathcal{C}\big(\mathcal{Z}(\mathbf{L})*\mathcal{Z}(\mathfrak{h})\big)\Big),$
@@ 2.3. Implicit Neural Representation
where $\mathbf{F}_{\theta}$ is an MLP with parameters $\theta$. This network takes the spatial and angular coordinates as input and outputs the corresponding pixel color value. The variables $\mathbf{x}$ and $\mathbf{y}$ denote the illuminated pixel coordinates on the display. Before the input coordinates $p$ are fed into the MLP, we apply positional encoding (Mildenhall et al., 2021) to preserve high-frequency details:
@@ 3. Evaluation and Discussion
Our hardware uses a $0.5^{\circ}$ engineered diffuser (Edmund 35-860) with five $2.5\,mm\times 3\,mm$ apertures positioned 10mm10\,$m$ from the imaging sensor as depicted in Fig. 4. We experimentally determine this spacing to ensure spatial invariance and sharp caustic patterns, satisfying Eq. (2). Moreover, we place the apertures such that we maximize the angle of rays arriving from a pixel landing on an imaging sensor. This aperture placement maximizes ray angles from pixels to the sensor, though sensor dimensions limit the maximum supported angle. In our implementation, the limited imaging sensor dimensions dictate the locations of these apertures, posing a constrain on the maximum angle we can support with our aperture arrangement. We 3D print a housing maintains 30mm30\,$m$ between diffuser and screen, as demonstrated in left image Fig. 5. We capture a stack of PSF for each aperture using a white LED with a $100\,\mu m$ pinhole (Thorlabs P100HKb), then mount the prototype in front of the display and activate pixels sequentially. To account for LCD backlight bleeding  [^6] https://pixiogaming.com/blogs/latest/understanding-backlight-bleed-in-ips-panels-causes-and-solutions, we sample nine screen positions to capture the full spatial distribution (Supplementary Sec. 1).
@@
We divide lensless captures into $9\times 9$ sub-aperture images using a $54\times 70$ pixel sliding window. Later, we feed them into a MLP that consists of 3 fully connected layers with sinusoidal activation functions and 32 channels each. To improve generalization, we inject random noise into the coordinates from Sec. 2.3: display (std: $5\times 10^{-3}$), angular (std: $1\times 10^{-2}$), and subview (std: $1\times 10^{-3}$). We then apply positional encoding Eq. (8) with 1, 5, and 10 frequency levels for display, angular, and spatial coordinates, respectively (see Fig. 3 and Supplementary Sec. 2). We optimize using Adam optimizer with an initial learning rate of $0.001$ that decays linearly over $800$ epochs and clip gradient norms to $1.0$. We train the model on independent color channels to mitigate crosstalk, which takes approximately one hour on an NVIDIA RTX 2070 GPU, while inference takes $0.01$ seconds per 1080p frame.
@@
To assess pixel consistency, we measure intensity variations in reconstructed light fields across continuous horizontal and vertical incident angles. We sample nine display positions with five measurements at different apertures per position, train the MLP on this data, and estimate pixel values at unseen angles. Fig. 5 shows smooth predicted intensity changes, demonstrating the model’s angle-dependent reconstruction capability. Our architecture outperforms a vanilla MLP baseline with fully connected ReLU layers across all metrics (Tbl. 2).
@@
To benchmark our method, we follow the ISO standard (ISO, 2008): align a camera with the display in a dark room, display full-screen white stimuli, and capture images from $10^{\circ}$ to $16^{\circ}$ vertical incident angles. Fig. 6 shows our method reproduces the ISO intensity trend, validating its physical plausibility. We capture light field data from nine display positions without camera rotation or controlled lighting, reducing measurement time and simplifying view-dependent calibration while maintaining comparable accuracy.
@@
Methods
PSNR
$\uparrow$
SSIM $\uparrow$ MSSIM $\uparrow$
Train
Inference
Ours 19.54 0.9165 0.9549 1 0.01
Vanilla 9.93 0.5327 0.8049 0.8 0.003
@@ Methods
Our $46.6^{\circ}$ angular coverage remains well below the $240^{\circ}$ required in professional display calibration. We could extend this range by co-optimizing the apertures and the diffuser design. Beyond angular coverage, pixel-wise training hinders scalability for full-panel characterization. To overcome these constraints, we outline several promising research directions. First, developing an end-to-end pipeline that eliminates PSF convolutions, cropping, and padding. Second, hash encoding (Müller et al., 2022) or Gaussian splatting could model light fields more efficiently.